OpenStudy (anonymous):

please help fan + medals!!! Sarah has been running a dog-walking business since 2010. She walks dogs twice a day, takes them to the park, and returns them to their homes. Each year, she has increased her fee by the same amount. The table shows what Sarah charged each customer for two given years of her business: Year Annual Dog-walking Fee 2010 $350 2014 $750 Q: What is the rate of change and initial value for Sarah’s business? How do you know?

Mehek (mehek14):

rate of change \(\dfrac{750-350}{2014-2010} = \dfrac{400}{4}=100\)

OpenStudy (rainbow_rocks03):

What do you think the answer is @TheHangingTree

Mehek (mehek14):

to find initial value 2010 = 100(350)+b

OpenStudy (rainbow_rocks03):

No direct answers plz.

OpenStudy (rainbow_rocks03):

It is against code violation

Mehek (mehek14):

100*350 = 3500

Mehek (mehek14):

@rainbow_rocks03 i've been on this site for a long time and I'm explaining so chill it

Mehek (mehek14):

2010 = 3500 + b

OpenStudy (anonymous):

I keep forgeting to subtract the topones, i always add or multiply them...

OpenStudy (anonymous):

top ones*

Mehek (mehek14):

2010-3500 = -1490 b = 1490

Mehek (mehek14):

so roc = 100 iv = -1490

OpenStudy (anonymous):

roc and iv?

OpenStudy (rainbow_rocks03):

You don't have to be rude @Mehek14

OpenStudy (rainbow_rocks03):

I was just saying

OpenStudy (anonymous):

all she said was to chill...

OpenStudy (anonymous):

@Mehek14 , @rainbow_rocks03 is right. you're supposed to guide the person to the answer and explain it as you go, bit just give the answer and explain why. I understand you intentions just remember that for next time.

Mehek (mehek14):

@Jennjuniper look at all the things I explained are you blind as well?

OpenStudy (anonymous):

i do have pretty bad eyes, but no. i am not blind. and i see what you were doing. i'm just letting you know a better way to do it for next time. This is a place to get homework HELP not homework answers. that's why these rules are in place.

Mehek (mehek14):

@Jennjuniper lol you're giving me advice? have you seen my ss? 91 which means ive been on this site longer than you meaning i know the rules more in depth than you and I helped not just directly gave the answer

OpenStudy (shaik0124):

Sahi jawab @Mehek14